OpenStudy (anonymous):

Can someone check if im correct?(medal) Use mathematical induction to prove the statement is true for all positive integers n. 10 + 30 + 60 + ... + 10n = 5n(n + 1)

OpenStudy (anonymous):

n=1 5(1)(1+1)=5(2)=10, which is the first term so its true 10=10 n=k 5k(k+1), assume that this statement is true n=k+1 5(k+1)[(k+1)+1]= 5k+5(k+2)=5k+5k+10=10k+10=5(k+1) which is equal to 5n(n+1), so its true i dont know if im right...

OpenStudy (anonymous):

nevermind :)

zepdrix (zepdrix):

oh figured it out? c: ok